### v1

# A Cross-Scene Benchmark for Open-World Drone Active Tracking

Abstract
Drone Visual Active Tracking aims to autonomously follow a target object by controlling the motion system based on visual observations, providing a more practical solution for effective tracking in dynamic environments. However, accurate Drone Visual Active Tracking using reinforcement learning remains challenging due to the absence of a unified benchmark, the complexity of open-world environments with frequent interference, and the diverse motion behavior of dynamic targets. To address these issues, we propose a unified cross-scene cross-domain benchmark for open-world drone active tracking called DAT. The DAT benchmark provides 24 visually complex environments to assess the algorithms’ cross-scene and cross-domain generalization abilities, and high-fidelity modeling of realistic robot dynamics. Additionally, we propose a reinforcement learning-based drone tracking method called R-VAT, which aims to improve the performance of drone tracking targets in complex scenarios. Specifically, inspired by curriculum learning, we introduce a Curriculum-Based Training strategy that progressively enhances the agent tracking performance in vast environments with complex interference. We design a goal-centered reward function to provide precise feedback to the drone agent, preventing targets farther from the center of view from receiving higher rewards than closer ones. This allows the drone to adapt to the diverse motion behavior of open-world targets. Experiments demonstrate that the R-VAT has about 400% improvement over the SOTA method in terms of the cumulative reward metric.
Figure 1: Example scenes of DAT benchmark. We construct 6 environments with 4 weather conditions for cross-scene and cross-domain generalization tasks, providing a unified benchmark for the VAT.
1 Introduction
Visual Active Tracking (VAT) aims to autonomously follow a target object by controlling the motion system of the tracker based on visual observations [73, 68]. It is widely used in real-world applications such as drone target tracking and security surveillance [20, 66, 70, 48]. Unlike visual passive tracking [4, 67, 28, 6, 9, 75, 60, 51], which involves proposing a 2D bounding box for the target on a frame-by-frame with a fixed camera pose, VAT actively adjusts the camera position to maintain the target within the field of view. Passive visual tracking often falls short in real-world scenarios due to the highly dynamic nature of most targets. Thus, VAT offers a more practical yet challenging solution for effective tracking in dynamic environments.
AD-VAT+ [73] D-VAT [18] AOT [33] DAT
Scenes 8 4 2 24
Targets 1 1 1 24
Sensors 2 3 2 6
Tracker Ground Drone Ground Both
Dynamics
Randomness
Table 1: Comparison of DAT benchmark with simulators where existing methods are located.
Recently, VAT methods have evolved into two main categories: pipeline VAT methods [34, 40, 14] and reinforcement learning-based VAT methods [18, 33, 17, 73]. Pipeline VAT methods employ a sequential framework where the visual tracking [29, 5, 56, 28] and control models are connected in series. Here, the object tracking model first processes the input image to estimate the target position and then the control model to generate the necessary control signals. While this modular design allows for clear task separation, these methods often require significant manual effort to label the training data, and the combination of modules requires extra effort for tuning and implementation. To address these issues, reinforcement learning-based VAT methods integrate visual tracking and control within a unified framework. These methods eliminate the need for separate tuning of the visual tracking and control modules by using a unified framework to map raw visual inputs directly to control actions. Therefore, the reinforcement learning-based VAT methods simplify system design and increase the efficiency of learning adaptive tracking behaviors in dynamic environments.
Unfortunately, achieving accurate drone visual active tracking with reinforcement learning remains challenging, partly for the following reasons. 1) Missing unified benchmark. Existing benchmark scenes are limited in scope, low in complexity, and few in number, and unable to adequately validate the performance of agents (see Tab. 1). Previous methods provide limited tracking targets and sensors, insufficient to establish benchmarks for different tasks. 2) Vast environments with complex interference. Open-world tracking involves large, dynamic environments with frequent interference. It brings significant challenges for agents to accurately follow targets. Training directly in these conditions often leads to slow convergence or difficulty in building robust tracking behaviors. 3) Complex targets with diverse motion behaviors. Open-world targets often exhibit complex, unpredictable behaviors, requiring agents to adapt to varied movements and orientations. Existing methods assume a fixed forward-facing view, leading to reward functions that inaccurately reflect tracking performance across different behaviors and perspectives.
Figure 2: A pipeline for drone visual active tracking.
To address the above limitations, we first propose a unified cross-scene benchmark for open-world drone active tracking (called DAT) that simulates the diversity and complexity of the real world as possible (see Fig. 1). Specifically, the DAT benchmark provides 24 visually complex scenes to validate the algorithms’ cross-scene and cross-domain generalization abilities. It offers comprehensive support for diverse tracking scenarios with 2 tracker types, 24 target types, and 6 different sensor types, with plug-and-play interfaces that facilitate the integration of custom robot models and controllers. To better replicate real-world conditions, DAT employs the webots simulation software [31] for high-fidelity modeling of realistic robot dynamics. It incorporates the Simulation of Urban Mobility (SUMO) [30] for managing target behavior, enabling diverse and efficient path modeling that surpasses traditional rule-based approaches. Second, we propose a novel drone visual active tracking with reinforcement learning method (called R-VAT), aiming to improve the performance of drone tracking targets in complex scenarios. Specifically, inspired by curriculum learning [58, 47, 35, 76], we propose a Curriculum-Based Training strategy that progressively improves agent performance in vast environments with complex interference. In addition, we design a goal-centered reward function to provide accurate feedback to the agent, enabling it to adapt to the target’s diverse motion behaviors. Unlike existing methods, this function is designed at the image level to prevent targets farther from the center of view from receiving higher rewards than closer ones.
Figure 3: Statistics and simulator component examples of the DAT. (a) Statistics on the 7 complexity aspects in DAT scenes. (b) Examples of challenging situations in various scenes. (c) Diagrams of the provided sensors. (d) Examples of available tracking targets.
To summarize, our contributions are as follows:
  • A comprehensive drone active tracking benchmark. We present the DAT benchmark, simulating 24 visually diverse, real-world environments with various trackers and targets, enabling rigorous cross-domain and cross-scene validation of algorithms.
  • A novel drone active tracking method. We propose the R-VAT, which leverages a Curriculum-Based Training strategy for tracking and a goal-centered reward function to enhance drone tracking performance in complex and dynamic environments.
  • Extensive experimental validation. Experiments validate the DAT benchmark usability, with R-VAT achieving about 400% improvement over the SOTA method in terms of the cumulative reward metric.
2 Task Definition of Drone Active Tracking
The DAT task seeks to train a model to control a drone for active target tracking in dynamic environments (see Fig. 2). Using visual and motion sensor data, the model learns actions to keep the target centered in view, ensuring robust performance across diverse scenarios.
Observation spaces. The target is initially positioned at the center of the camera’s field of view, and the observation space comprises data acquired from sensors (e.g., RGB images with a resolution of $84\times 84$).
Action spaces. There are two options for the action space: discrete action space and continuous action space. The discrete action space is a 7-dimensional vector, representing movements such as forward, backward, left translation, right translation, counterclockwise rotation, clockwise rotation, and stopping. The speed corresponding to each action must be set by the user. The continuous action space is a 4-dimensional vector that defines speed values in the forward, lateral, vertical, and yaw directions.
Success criterion of DAT task. We define a success criterion when the model can keep the target object, which is initially located at the center of the camera’s field of view, in the middle of the image for a long duration.
Challenges. Open-world drone active tracking presents substantial challenges primarily due to the scarcity of real-world data and the high costs and risks associated with real-world trial-and-error. This highlights the critical need for building complex and diverse simulation environments. Additionally, open-world scenes are characterized by high diversity and dynamic elements, which introduce complex interferences, further challenging the model’s robustness. Moreover, real-world tracking targets exhibit diverse and unpredictable behaviors, making it difficult for models to adapt. Therefore, improving the adaptability of models to track complex targets effectively is another key challenge.
3 DAT Benchmark with Diverse Settings
We develop the DAT benchmark (see Fig. 3) to evaluate the model’s cross-scene and cross-domain adaptability in 6 large outdoor environments with 4 weather conditions.
3.1 Diverse Scene Construction
The construction of the DAT scene aims to simulate the diversity and complexity of the real world. To enhance diversity, we select 6 categories of outdoor scenes and 4 weather conditions, covering almost all natural conditions that applications may face. To simulate the real-world complexity, we model 7 aspects: scene area, building density, color richness, road density, terrain density, tree density, and tunnel density. Specifically, the scene area refers to the extent of the scene, building density is the ratio of the number of buildings to the scene area, and color richness is the number of dominant colors in the scene. These three aspects primarily depict the complexity of the visual background. Road density is measured by evaluating the density of complex elements such as intersections, terrain density is determined by assessing the density of special terrain features like mountains in the scene. These two aspects primarily depict the complexity of tracking target behavior. The tree density and tunnel density are calculated as the ratio of the number of trees and tunnels to the scene area, and are used to measure the level of visual occlusion within the scene.
As shown in Fig. 3(a), the six different scenarios exhibit unique and realistic complexity across the seven aspects:
•  Citystreet scene covers an area of 0.7 square kilometers. It has a road density of 38.7 and a tree density of 97.5, mainly testing the agent’s efficiency against tree occlusions.
•  Village scene spans 1.4 square kilometers. This scene features a mountainous terrain density of 20.1 and a tunnel density of 6.3, requiring the agent to predict the target’s movement when it is fully obscured by tunnels.
•  Downtown scene covers 0.8 square kilometers. The scene includes complex road elements and a very high building density of 304.9, challenging the agent’s tracking accuracy and obstacle avoidance capabilities.
•  Lake scene encompasses 1.6 square kilometers. The density of road elements is 68.1, and the richness of background colors is rated at 5, challenging the agent’s robustness across varying features and colors.
•  Farmland scene covers an area of 0.7 square kilometers. The color richness is rated at 5, with multiple color patches spread throughout the scene, posing a significant challenge to the agent’s adaptability to multi-color environments.
•  Desert scene covers 1.1 square kilometers. It includes a mountainous terrain density of 37.1 and a road element density of 31.0. Some roads are covered by sand, testing the agent’s adaptability to such challenging conditions.
The four weather conditions mainly test the agent’s cross-domain adaptability. Foggy reduces visibility, with a visibility range of 400m to 2000m. Night reduces brightness and light uniformity, and snow changes the color tone of the environment. The above 24 scenes can fully measure the performance of the agent active adaptation (see Fig. 1).
3.2 Various Trackers and Targets Construction
Drone Active Tracking in real-world applications involves various trackers and targets depending on the specific task. For example, drones are typically used as trackers, with automobiles as targets in security monitoring. The DAT benchmark provides diverse trackers and targets, making it adaptable for different use cases.
Tracker. The DAT benchmark supports two tracker types: drones and ground robots. The drone used is the DJI Matrice 100 [12], with a default flight controller [53]. The DAT benchmark provides visual and motion sensors (see Fig. 3(c)) that can obtain the state parameters of the drone relative to the world coordinate system. The drone position and velocity are obtained via GPS, angular velocity via gyroscope, Euler angles are obtained from the IMU and can be converted into quaternion, and acceleration via accelerometer. The LiDAR provides point cloud data. To better replicate real-world conditions, drones are equipped with a 3-axis gimbal, similar to those on commercial drones like the DJI Mini 3 Pro [13], enabling precise camera adjustments and adding complexity beyond static point trackers. For ground robots, users can select from various robot models. The benchmark employs webots simulation software to provide a high-fidelity dynamic engine, accurately modeling tracker movements, collisions, and interactions, crucial for evaluating tracking algorithms in challenging scenarios. Additionally, we provide common parameters for further task customization. See Appendix B for details.
Targets. The DAT benchmark includes five categories of targets: automobile, motorbike, pedestrian, wheel robot, and legged robot, providing a total of 24 tracking targets, each with built-in controllers (see Fig. 3(d)). To support custom robot designs, a plug-and-play interface is available, allowing users to easily integrate robot models and controllers into the benchmark. See Appendix B for details.
Target Management. Realistic target behavior and path diversity are essential for simulating authentic environments. The DAT benchmark uses the SUMO traffic simulator to manage all scene targets. For example, SUMO generates random vehicle paths and dynamically controls actions like acceleration, lane changes, and stops. It can also refresh vehicles based on configurations and randomize their types and colors, creating a high-fidelity traffic system.
4 VAT with Reinforcement Learning
In this paper, we primarily focus on visual active tracking (VAT), a core task within the DAT benchmark. We propose a drone visual active tracking with reinforcement learning method (called R-VAT), aiming to improve the performance of tracking targets in complex scenes. As shown in Fig. 2, we model drone active tracking as a Markov Decision Process (MDP) and train a Drone Agent to track a target in the open scene. At time step $t$, the Drone Agent takes the current state $s_{t}$ as input and selects an action. Then the reward is calculated via Eq. 2. The trajectories are collected to train the agent via the Curriculum-Based Training strategy.
4.1 MDP for Drone Active Tracking
[…]
State $\boldsymbol{\mathcal{S}}$ is the visual information of the scene. At each time step $t$, the camera captures one image of size $84\times 84$ as the current state.
Action $\boldsymbol{\mathcal{A}}$ is a set of discrete actions that the drone can take, including forward, backward, left translation, right translation, counterclockwise rotation, clockwise rotation, and stopping, respectively. At each time step, the Drone Agent selects an action $a_{t}\in\boldsymbol{\mathcal{A}}$ based on the state $s_{t}$ and actively controls the camera movement.
Transition $\mathcal{T}(s_{t},a_{t})$ is a function $\boldsymbol{\mathcal{T}}:\boldsymbol{\mathcal{S}}\times\boldsymbol{\mathcal{A}}\rightarrow\boldsymbol{\mathcal{S}}$ that maps $s_{t}$ to $s_{t+1}$. In this paper, we use the webots dynamics engine to provide a realistic transition function.
Reward $\mathcal{R}(s_{t},a_{t})$ is the reward function. The reward we designed is constructed only at the image level. The details are given in Sec. 4.3.
Network structure of Drone Agent. Since the VAT task requires the agent to maintain long-term tracking, a dynamics model is essential for capturing temporal information. Therefore, similar to previous works [33, 73, 17], we select a CNN combined with a gated recurrent unit [11] network architecture as the backbone (see Appendix C).
4.2 Curriculum Learning for Agent Training
The DAT environment contains numerous dynamic targets and diverse obstacles, making it challenging for the agent to discover sufficient successful trajectories. This difficulty results in low convergence rates and limited performance. Progressively training the agent from simpler to more complex environments increases the percentage of successful trajectories, enhancing performance and accelerating learning for the final task [57]. Therefore, we propose a Curriculum-Based Training (CBT) strategy for tracking to optimize agent training in complex environments.
Training Algorithm. To address the challenges of learning in complex environments, we employ the Proximal Policy Optimization (PPO) [49] algorithm, which is widely recognized for its robustness and efficiency in complex continuous control tasks. Given the complexity of the environment, we adopt a multi-stage training strategy, which divides the model training into two stages to build the agent’s tracking capabilities progressively. The first stage $\mathcal{E}_{1}$, consists of a simplified environment where the target trajectory is a straight line, and there are no obstacles or visual interferences. The agent learns to keep the target centered in the image through the reward signal from Eq. 2. Once the reward obtained for agent in first stage $\mathcal{E}_{1}$ converges as:
$\small\frac{1}{T}\sum_{t=1}^{T}r_{t}\geq\eta,$ (1)
where $r_{t}$ represents the reward obtained by the agent at time step $t$, $T$ is the total number of time steps over which the reward is averaged, and $\eta$ is a predefined threshold indicating satisfactory performance. Once this condition is satisfied, the agent transitions to the second stage $\mathcal{E}_{2}$, where the agent encounters more varied tracking target movements and more complex visual information, such as tree occlusions, crosswalks, and poles. The goal for the agent is to develop stronger task generalization abilities based on task understanding gained in the first stage.
Data Augmentation. While simpler settings facilitate the agent’s learning of task objectives, they also heighten the risk of the agent rapidly converging to a locally optimal action distribution, which can undermine the exploration process. Consequently, implementing data augmentation is essential for enhancing the agent’s exploratory capabilities. This is achieved through the randomization of the drone’s initial position and orientation relative to the target, which necessitates a broader range of actions from the agent to maximize rewards. Moreover, to enhance the agent’s spatial perception ability, randomization is also introduced in its gimbal pitch angle. See Appendix C for details.
Figure 4: Schematic diagram of reward acquisition.
4.3 Goal-Centered Reward Design
Open-world targets often exhibit complex, unpredictable behaviors, requiring agents to adapt to varied movements and orientations. However, existing methods assume the drone is parallel to the target, leading to reward functions that inaccurately reflect tracking performance across different behaviors and perspectives (see Appendix C for theoretical proof). To address this, our reward function is defined only at the image level, relying on the target position within the image $I\in\mathbb{R}^{84\times 84}$. The reward decreases as the target moves away from the center $I_{O}$ of the image, and no reward is given if the target is outside as:
$\small r_{t}=\left\{\begin{array}[]{lr}\tanh{(\alpha(1-x)^{3})}\ \ I_{target}\in I_{clip}\\ 0\ \ otherwise\end{array},\right.$ (2)
where $x=\frac{\mid p_{vc}-I_{OG}\mid}{\mid E_{G_{c}}-I_{OG}\mid}$ represents the position of the target relative to the center of the image, which is calculated by projecting the image onto the ground. Specifically, we calculate the ratio of the distance from the projection $I_{OG}$ of the image center on the ground to the center $p_{vc}$ of the target (line segment $\mid\!p_{vc}\!-\!I_{OG}\!\!\mid$) and the distance from the intersection $E_{G_{c}}$ of the line connecting these two centers with the edge of the image projection to the center $I_{OG}$ (line segment $\mid\!E_{G_{c}}\!-\!I_{OG}\!\!\mid$), as shown in Fig. 4(b). The attenuation degree of $r_{t}$ can be adjusted using the hyperparameter $\alpha$, set to 4. The $\tanh(\cdot)$ provides a strong indication of the task target due to its relatively quick decay at the center of the image. $I_{clip}$ is the truncated image range set to prevent the drone from keeping the target at the edge of the image for a long time. The truncation of the image can be controlled using the hyperparameter $\lambda_{clip}$ as:
$\small\lambda_{clip}=\frac{W_{I_{clip}}}{W},$ (3)
where $W$ and $W_{I_{clip}}$ are the widths of the original image and the truncated image, respectively, we set $\lambda_{clip}=0.7$.
The reward function relies on the judgment of the corner points of the image: Left Up (LU), Left Down (LD), Right Up (RU), and Right Down (RD) and the target coordinates along with their projections. To map corner points onto the ground, the effective focal length of the camera ($f$) must first be estimated, as shown in Fig. 4(a). This can be calculated using the pinhole imaging principle [8], as follows:
$\small f=\frac{W}{2\tan{(\frac{1}{2}FoV)}},$ (4)
where $W$ is the camera image width (in pixels) and $FoV$ is the field of view of the tracker camera. After obtaining the focal length, we project the image corner points onto the target plane to check if the target falls within the image range, as shown in Fig. 4(a). This involves finding intersections between lines extending from the camera optical center through its corner points and the ground plane.
[…]
$\small\left\{\begin{array}[]{lr}l_{LUC}:\frac{x}{-f}=\frac{2y}{-W}=\frac{2z}{H}\\[4.5pt] l_{LDC}:\frac{x}{-f}=\frac{2y}{-W}=\frac{2z}{-H}\\[4.5pt] l_{RUC}:\frac{x}{-f}=\frac{2y}{W}=\frac{2z}{H}\\[4.5pt] l_{RDC}:\frac{x}{-f}=\frac{2y}{W}=\frac{2z}{-H}\\[4.5pt] \end{array}\right.,$ (5)
[…]
In coordinate transformations, a homogeneous transformation matrix (HTM) can be used to map a plane into another coordinate system. Suppose the analytical equation of the original plane $P_{0}$ is given by $A_{0}x+B_{0}y+C_{0}z+D_{0}\!=\!0$. The transformed plane $P_{1}$ has the equation $A_{1}x\!+B_{1}y\!+C_{1}z\!+D_{1}\!=\!0$. The homogeneous vectors for the two planes are $p_{0}\!\!=\![\hat{n_{0}},\!D_{0}]$ and $p_{1}\!\!=\![\hat{n_{1}},\!D_{1}]$, where $\hat{n_{0}}$ and $\hat{n_{1}}$ are the normal vectors. The HTM $T_{01}$ from plane $P_{0}$ to $P_{1}$ as:
$\small T_{01}=\begin{bmatrix}R_{01}&t_{01}\\ 0&1\\ \end{bmatrix},$ (6)
where $R_{01}$ and $t_{01}$ represent the rotation matrix and translation vector, respectively. The normal vector $\hat{n_{1}}$ of $P_{1}$ can be expressed as: $\hat{n_{1}}^{T}\!\!=\!\!R_{01}\hat{n_{0}}^{T}$.
Assume that the points $\mathbf{p_{0}}=[x_{0},y_{0},z_{0}]^{T}$ and $\mathbf{p_{1}}=[x_{1},y_{1},z_{1}]^{T}$ lie on the planes $P_{0}$ and $P_{1}$, respectively. These two points satisfy as follows:
$\hat{n_{0}}\mathbf{p_{0}}+D_{0}$ $=0,$ (7)
$\hat{n_{1}}\mathbf{p_{1}}+D_{1}$ $=0.$
Based on coordinate transformation principles, the relationship between the two points can be derived as $\mathbf{p_{1}}=R_{ct}\mathbf{p_{0}}+t_{ct}$. Therefore, by solving the simultaneous equations, the expression for constant term $D_{1}$ can be obtained:
$\small D_{1}=D_{0}-\hat{n_{1}}t_{01}.$ (8)
[…]
$\small A_{G_{c}}x+B_{G_{c}}y+C_{G_{c}}z+D_{G_{c}}=0.$ (9)
[…]
where $LU_{G}$, $LD_{G}$, $RU_{G}$ and $RD_{G}$ represent the mapping points of $LU$, $LD$, $RU$ and $RD$ on the ground plane. The expressions of $t_{LU}$, $t_{LD}$, $t_{RU}$, and $t_{RD}$ are as follows:
$\small\left\{\begin{array}[]{lr}t_{LU}=D_{G_{c}}(A_{G_{c}}f+\frac{1}{2}B_{G_{c}}W-\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] t_{LD}=D_{G_{c}}(A_{G_{c}}f+\frac{1}{2}B_{G_{c}}W+\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] t_{RU}=D_{G_{c}}(A_{G_{c}}f-\frac{1}{2}B_{G_{c}}W-\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] t_{RD}=D_{G_{c}}(A_{G_{c}}f-\frac{1}{2}B_{G_{c}}W+\frac{1}{2}C_{G_{c}}H)^{-1}\\[4.5pt] \end{array}\right..$ (11)
Method citystreet desert village downtown lake farmland
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
Within Scene
AOT $49_{\pm 3}$ $0.25_{\pm 0.02}$ $9_{\pm 1}$ $0.06_{\pm 0.00}$ $46_{\pm 5}$ $0.23_{\pm 0.03}$ $54_{\pm 5}$ $0.29_{\pm 0.01}$ $47_{\pm 3}$ $0.24_{\pm 0.02}$ $60_{\pm 25}$ $0.23_{\pm 0.01}$
D-VAT $48_{\pm 8}$ $0.26_{\pm 0.02}$ $47_{\pm 13}$ $0.26_{\pm 0.04}$ $44_{\pm 8}$ $0.22_{\pm 0.05}$ $9_{\pm 1}$ $0.06_{\pm 0.01}$ $46_{\pm 8}$ $0.26_{\pm 0.06}$ $13_{\pm 1}$ $0.07_{\pm 0.00}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{0.80}_{\pm 0.30}$ $\textbf{307}_{\pm 124}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{239}_{\pm 134}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{203}_{\pm 119}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{181}_{\pm 116}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$
Cross Scene
AOT $48_{\pm 5}$ $0.24_{\pm 0.02}$ $9_{\pm 1}$ $0.06_{\pm 0.00}$ $52_{\pm 11}$ $0.25_{\pm 0.03}$ $52_{\pm 6}$ $0.28_{\pm 0.03}$ $48_{\pm 5}$ $0.24_{\pm 0.02}$ $49_{\pm 7}$ $0.24_{\pm 0.02}$
D-VAT $49_{\pm 9}$ $0.26_{\pm 0.04}$ $48_{\pm 8}$ $0.27_{\pm 0.03}$ $50_{\pm 14}$ $0.24_{\pm 0.06}$ $8_{\pm 1}$ $0.05_{\pm 0.00}$ $51_{\pm 14}$ $0.25_{\pm 0.06}$ $13_{\pm 1}$ $0.07_{\pm 0.01}$
R-VAT $\textbf{144}_{\pm 111}$ $\textbf{0.52}_{\pm 0.29}$ $\textbf{230}_{\pm 115}$ $\textbf{0.67}_{\pm 0.27}$ $\textbf{156}_{\pm 119}$ $\textbf{0.55}_{\pm 0.31}$ $\textbf{201}_{\pm 121}$ $\textbf{0.64}_{\pm 0.30}$ $\textbf{163}_{\pm 115}$ $\textbf{0.51}_{\pm 0.29}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$
Table 2: Results of within and across scenes on the DAT benchmark.
[…]
Therefore, the coordinates of the car $p_{vc}$, along with the four intersection points ($LU_{G}$, $LD_{G}$, $RU_{G}$, and $RD_{G}$) and the image center $I_{OG}$ can be utilized to calculate the reward.
5 Experiments
Figure 5: Reward values during the training phase of our method and SOTA methods across six scenes.
5.1 Experimental Settings
[…]
Metrics. We use cumulative reward ($CR=\sum_{t=1}^{E_{l}}{r_{ct}}$) and tracking success rate ($TSR=\frac{1}{E_{ml}}\sum_{t=1}^{E_{l}}{r_{dt}}\times 100\%$) to evaluate the agent performance. The $CR$ primarily measures the tracker’s ability to keep the target centered in the image, where $r_{ct}$ is the dense reward at time step $t$ from Eq. 2, and $E_{l}$ is the length of the episode. The $TSR$ measures the tracker’s ability to maintain the target within the field of view, where $r_{dt}$ is the sparse reward (See Appendix C) at time step $t$ and $r_{dt}=1$ means the target within the field of view, $E_{ml}$ is the maximum episode length. Additionally, we position each agent relative to the target at $[0,\frac{\pi}{2},\pi,\frac{3\pi}{2}]~{}\text{rad}$ for 10 episodes each, totally 40 episodes. The mean and variance of these results are then calculated as evaluation metrics for the specific map. The model’s final cross-scene and cross-domain performance are both obtained by averaging results across different scenes.
Baselines. We reproduce the two SOTA VAT methods: AOT [33] and D-VAT [18]. See Appendix D for details.
Method night foggy snow
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
AOT $42_{\pm 4}$ $0.22_{\pm 0.02}$ $44_{\pm 7}$ $0.22_{\pm 0.02}$ $44_{\pm 7}$ $0.22_{\pm 0.02}$
D-VAT $35_{\pm 7}$ $0.19_{\pm 0.03}$ $37_{\pm 7}$ $0.19_{\pm 0.03}$ $34_{\pm 6}$ $0.19_{\pm 0.03}$
R-VAT $\textbf{217}_{\pm 125}$ $\textbf{0.64}_{\pm 0.32}$ $\textbf{243}_{\pm 114}$ $\textbf{0.76}_{\pm 0.26}$ $\textbf{178}_{\pm 105}$ $\textbf{0.60}_{\pm 0.26}$
Table 3: Results of cross domain on the DAT benchmark.
5.2 Comparison Experiments
Version citystreet desert village downtown lake farmland
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$ CR $TSR$
Within Scene
w/o CBT $54_{\pm 7}$ $0.30_{\pm 0.05}$ $302_{\pm 99}$ $0.75_{\pm 0.32}$ $\textbf{255}_{\pm 118}$ $\textbf{0.82}_{\pm 0.16}$ $57_{\pm 15}$ $0.33_{\pm 0.06}$ $\textbf{187}_{\pm 123}$ $0.57_{\pm 0.28}$ $46_{\pm 2}$ $0.23_{\pm 0.01}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{0.80}_{\pm 0.30}$ $\textbf{307}_{\pm 124}$ $\textbf{0.84}_{\pm 0.29}$ $239_{\pm 134}$ $0.73_{\pm 0.32}$ $\textbf{203}_{\pm 119}$ $\textbf{0.65}_{\pm 0.30}$ $181_{\pm 116}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$
Cross Scene
w/o CBT $39_{\pm 12}$ $0.23_{\pm 0.08}$ $\textbf{243}_{\pm 112}$ $0.69_{\pm 0.28}$ $\textbf{159}_{\pm 108}$ $0.52_{\pm 0.26}$ $50_{\pm 9}$ $0.28_{\pm 0.03}$ $139_{\pm 88}$ $0.49_{\pm 0.26}$ $53_{\pm 16}$ $0.26_{\pm 0.07}$
R-VAT $\textbf{144}_{\pm 111}$ $\textbf{0.52}_{\pm 0.29}$ $230_{\pm 115}$ $\textbf{0.67}_{\pm 0.27}$ $156_{\pm 119}$ $\textbf{0.55}_{\pm 0.31}$ $\textbf{201}_{\pm 121}$ $\textbf{0.64}_{\pm 0.30}$ $\textbf{163}_{\pm 115}$ $\textbf{0.51}_{\pm 0.29}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$
Table 4: Effectiveness of Curriculum-Based Training strategy within and across scenes on the DAT benchmark.
We compare our R-VAT with the SOTA methods for within-scene performance and cross-scene cross-domain generalization performance on the DAT benchmark. As shown in Fig. 5, our method achieves consistently higher and steadily increasing rewards throughout training, demonstrating its effectiveness in improving tracking performance across scenarios compared to other methods.
Within-scene performance. We train the model on all scenes and evaluate it on the original scene. Our R-VAT performs significantly better than other methods as shown in Tab. 2. For the $CR$, the average performance improvement on six maps relative to the D-VAT method is $591\%(35\!\rightarrow\!242)$. Regarding the $TSR$, the average enhancement is $279\%(0.19\!\rightarrow\!0.72)$.
Cross-scene performance. Our method demonstrates strong cross-scene generalization, as shown in Tab. 2. Specifically, R-VAT achieves a $376\%(37\!\rightarrow\!176)$ improvement in average $CR$ and a $200\%(0.19\!\rightarrow\!0.57)$ improvement in average $TSR$ compared to D-VAT.
Cross-domain performance. As shown in Tab. 3, our method outperforms existing methods significantly in cross-domain generalization. Specifically, the R-VAT demonstrates an average $CR$ enhancement of $509\%(35\!\rightarrow\!213)$ relative to D-VAT and $TSR$ boost of $253\%(0.19\!\rightarrow\!0.67)$.
Version night foggy snow
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
w/o CBT $136_{\pm 58}$ $0.49_{\pm 0.16}$ $138_{\pm 65}$ $0.48_{\pm 0.17}$ $107_{\pm 41}$ $0.39_{\pm 0.14}$
R-VAT $\textbf{217}_{\pm 125}$ $\textbf{0.64}_{\pm 0.32}$ $\textbf{243}_{\pm 114}$ $\textbf{0.76}_{\pm 0.26}$ $\textbf{178}_{\pm 105}$ $\textbf{0.60}_{\pm 0.26}$
Table 5: Effectiveness of Curriculum-Based Training strategy across domain on the DAT benchmark.
5.3 Ablation Experiments
We conduct ablation experiments on a Curriculum-Based Training strategy to demonstrate its effectiveness in challenging scenarios. Moreover, we conduct a validation on the farmland map to assess the effectiveness of the data augmentation techniques discussed in Sec. 4.2, as well as the reward design outlined in Sec. 4.3. Detailed descriptions of these methods are provided in Appendix E.
Effectiveness of Curriculum-Based Training strategy. As shown in Tab. 4 and Tab. 5, the R-VAT method exhibits the most significant improvements over the reinforcement learning approach on maps citystreet, downtown, and farmland, which are characterized by dense elements and visual complexity. In the within-scene, cross-scene, and cross-domain tests, enhancements in the $CR$ are $61\%(150\!\rightarrow\!242)$, $54\%(114\!\rightarrow\!176)$, and $68\%(127\!\rightarrow\!213)$, respectively, while improvements in the $TSR$ are $44\%(0.50\!\rightarrow\!0.72)$, $44\%(0.41\!\rightarrow\!0.59)$, and $49\%(0.45\!\rightarrow\!0.67)$.
Effectiveness of data augmentation. We apply randomization of angles (AR), horizontal displacement (HR), vertical displacement (VR), and gimbal pitch angle (PR) to enhance the tracking ability of the agent (see Sec. 4.2 for details). As shown in Tab. 6, adjusting AR or PR significantly improves the agent’s performance, indicating that enhancing exploratory behavior through angle variations and improving spatial awareness by randomizing the gimbal pitch angle yield positive outcomes.
Effectiveness of reward design. We contrast the performance of the R-VAT method when using the reward defined in Eq. 2 and that in [18]. As shown in Tab. 7, significant performance enhancements (about 800% improvement in $TSR$ across-scene and cross-domain) are evident on the farmland map with the utilization of Eq. 2, underscoring the pronounced effectiveness of the proposed reward design in this work. More experiments can be seen in Appendix E.
Version Within-Scene Cross-Scene Cross-Domain
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
w/o AR $106_{\pm 88}$ $0.44_{\pm 0.23}$ $92_{\pm 72}$ $0.37_{\pm 0.19}$ $80_{\pm 63}$ $0.36_{\pm 0.19}$
w/o HR $174_{\pm 118}$ $0.49_{\pm 0.30}$ $148_{\pm 129}$ $0.48_{\pm 0.32}$ $184_{\pm 124}$ $0.57_{\pm 0.30}$
w/o VR $211_{\pm 138}$ $0.63_{\pm 0.35}$ $161_{\pm 115}$ $0.54_{\pm 0.32}$ $203_{\pm 117}$ $0.60_{\pm 0.32}$
w/o PR $139_{\pm 119}$ $0.61_{\pm 0.33}$ $124_{\pm 85}$ $0.48_{\pm 0.25}$ $145_{\pm 122}$ $0.52_{\pm 0.28}$
R-VAT $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$ $\textbf{222}_{\pm 110}$ $\textbf{0.66}_{\pm 0.27}$
Table 6: Effectiveness of augmentation on the DAT benchmark.
Version Within-Scene Cross-Scene Cross-Domain
$CR$ $TSR$ $CR$ $TSR$ $CR$ $TSR$
$\text{R}_{\text{D-VAT}}$ ${9}_{\pm 1}$ ${0.06}_{\pm 0.00}$ ${9}_{\pm 1}$ ${0.06}_{\pm 0.01}$ ${9}_{\pm 0}$ ${0.06}_{\pm 0.00}$
R-VAT $\textbf{243}_{\pm 117}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{162}_{\pm 107}$ $\textbf{0.54}_{\pm 0.26}$ $\textbf{222}_{\pm 110}$ $\textbf{0.66}_{\pm 0.27}$
Table 7: Effectiveness of reward design on the DAT benchmark.
6 Conclusion and Potential Impacts
In this paper, we propose a unified cross-scene cross-domain benchmark for open-world drone active tracking, called DAT. The DAT benchmark provides 24 visually complex environments to assess the algorithms’ cross-scene and cross-domain generalization abilities, and high-fidelity modeling of realistic robot dynamics. Additionally, we propose a reinforcement learning-based drone tracking method called R-VAT, aiming to improve the performance of drone tracking targets in complex scenarios. Specifically, inspired by curriculum learning, we implement a Curriculum-Based Training strategy that systematically improves agent performance in increasingly challenging scenarios. We design a goal-centered reward function to deliver precise feedback to the drone agent, preventing targets farther from the center of view from receiving higher rewards than closer ones. This allows the drone agent to adapt to the diverse motion behavior of open-world targets. Experiments demonstrate that our method is significantly superior to the SOTA method.
The DAT benchmark is a high-fidelity, cross-scene cross-domain benchmark for advancing research in drone visual active tracking. It has the potential to impact several key areas, including: 1) Forgetting in Reinforcement Learning, 2) Robustness in Reinforcement Learning, 3) Multi-Agent Reinforcement Learning, and 4) Sim-to-Real Deployment.
Appendix A Related Work
A.1 Passive Object Tracking
Most of the proposed visual tracking benchmarks belong to passive visual tracking. LaSOT [21] and OTB2015 [64] benchmarks contain a large number of ground-based videos. These benchmarks include target videos, and the tracking algorithms utilize both the video frames and the target labels for tracking. However, ground cameras tend to be affected by occlusion and suffer from the shortcoming of limited perceptual range, so the need for drone viewpoint tracking is gradually increasing in practical applications. UAV123 [37] and VisDrone2019 [19] benchmarks are proposed for drone viewpoint, expanding the spatial dimension of perception. Meanwhile, the single-object tracking benchmarks have difficulties for many targets. MOT20 [16] and TAO [15] benchmarks are proposed for multi-object tracking to solve the above problems. In addition, the above benchmarks include videos from the RGB camera. The RGB camera’s recognition capabilities are limited in complex scenes, such as ocean environments, and challenging weather conditions, including nighttime and foggy. IPATCH [41] provides extra infrared images and other sensors like GPS to supplement the information of the sea scene. Huang et al. propose Anti-UAV410 [25], which provides infrared camera images for drone tracking.
Visual object tracking methods can be categorized into three main types: Tracking by Detection, Detection and Tracking (D&T), and pure tracking. Tracking by Detection methods [6, 62, 7] treat tracking as a sequence of independent detection tasks. These methods use object detection algorithms [44, 46] to identify the target object in each frame, connecting the detections through data association methods [27, 63] for continuous tracking. While effective in multi-target tracking, these methods may suffer from high computational demands and issues with target occlusion. D&T approaches [59, 72, 42] integrate detection and tracking, creating end-to-end models that ensure seamless information flow and reduce redundant calculations through shared feature extraction networks. Pure tracking methods can be categorized into two main types: Correlation Filters (CF) [65, 38, 24] and Siamese Networks (SN) [29, 5, 56]. CF-based models train correlation filters on regions of interest, while SN-based models compare target templates with search areas to enable precise single-target tracking.
A.2 Visual Active Tracking
Passive visual tracking often falls short in real-world scenarios due to the highly dynamic nature of most targets. Visual Active Tracking (VAT) aims to autonomously follow a target object by controlling the motion system of the tracker based on visual observations [34, 73, 68]. Thus, VAT offers a more practical yet challenging solution for effective tracking in dynamic environments. Maalouf et al. [34] propose a two-stage tracking method (named FAn), which is based on a tracking model and a PID control model. This method accomplishes the fusion of perception and decision-making by transferring control information from the visual tracking model to the control model. However, the visual network necessitates extensive human labeling effort and the control model requires parameter adjustments for each scene, significantly constraining the model’s generalizability. Recently, many approaches [33, 17, 73, 18] model the VAT task as a Markov Decision Process and employ end-to-end training with reinforcement learning, resulting in a significant enhancement of the agent’s generalizability.
The complexity and diversity of VAT benchmarks are crucial for training agents with high generalizability. One common approach [18, 17, 73] to enhancing environmental diversity involves modifying texture features and lighting conditions within a single scene. However, these methods often result in low scene fidelity and unrealistic object placement. While UE4 [22] is used to create photorealistic environments in some benchmarks [73, 33], these benchmarks still face limitations in diversity and map size. Furthermore, the scenarios provided by these methods are often task-specific, offering limited configurability and lacking a unified benchmark for VAT tasks.
Existing approaches to VAT frequently neglect the randomness of target trajectories and the scalability of platforms. Target trajectories are typically predefined by rule-based patterns [18, 17, 33], which significantly restrict the exploration space. Zhong et al. [73] introduce learnable agents as targets, increasing trajectory randomness but adding additional cost. Most benchmarks provide only a single category of target [18, 17, 73, 33], limiting scalability and necessitating repetitive work for environment development. Zhou et al. [74] utilize CoppeliaSim [3] to provide five categories of noncooperative space objects. However, the use of a solid black background makes it unsuitable for general VAT scenarios. In contrast, our environment supports diverse, real-world target types and offers unified, lightweight management of target behaviors, ensuring both rationality and randomness in their actions.
A.3 Reinforcement Learning in Visual Tracking
Reinforcement learning (RL) is commonly used in visual object tracking [69, 45, 71]. Song et al. [52] propose a decision-making mechanism based on hierarchical reinforcement learning (HRL), which achieves state-of-the-art performance while maintaining a balance between accuracy and computational efficiency. However, the actions generated by reinforcement learning in the aforementioned work cannot directly influence the camera’s viewpoint, thereby failing to fully leverage the decision-making capabilities. Real-world applications increasingly require robust tracking in highly dynamic scenes, motivating researchers to explore reinforcement learning agents for effectively synchronizing visual perception and decision-making in VAT tasks. Dionigi et al. [18] demonstrate the feasibility of reinforcement learning for drone VAT missions. However, the assumption of a fixed-forward perspective limits its applicability in real-world tasks.
A.4 Curriculum Learning in Robot Control
Curriculum Learning (CL) is a training strategy that mimics a human curriculum by training models on simpler subsets of data at first and gradually expanding to larger and more difficult subsets of data until they are trained on the entire dataset. As for robot control, reinforcement learning training is difficult due to the complexity of the training scenarios and the large action spaces. Therefore, curriculum learning is often required to reduce the difficulty of agent training. For instance, many works improve the walking ability of legged robots by adjusting terrain parameters through curriculum learning [47, 35]. Other studies improve the pushing and grasping performance of robotic arms by progressively increasing task difficulty [32, 55, 39].
In this paper, Curriculum Learning is introduced in the VAT task, and the training environment is transitioned from simple features to complex scenarios to achieve successful tracking of agent in complex outdoor environments.
Figure 6: Examples of DAT benchmark targets. (a) Illustration of tracking targets for 10 types of automobile and 2 types of motorbike. (b) Illustration of tracking targets for the pedestrian type. (c) Illustration of tracking targets for 5 types of wheeled robot. (d) Illustration of tracking targets for 6 types of legged robot.
Appendix B More Details of DAT Benchmark
Scenario Construction. In this study, three scenarios: citystreet, downtown, and lake are derived from real-world locations. The corresponding map segments for these scenarios are extracted from OpenStreetMap (OSM) [2], and road information, such as the number of lanes and intersection rules, is edited using the JOSM tool [1]. Then configurations are then converted into webots-compatible assets. Specifically, the citystreet scenario is based on a small town in Los Angeles, the downtown scenario is derived from Manhattan, and the lake scenario is modeled after Wolf Lake Memorial Park in Indiana. In contrast, the village, desert, and farmland maps possess complex and unique features that are not adequately captured by OpenStreetMap (OSM) data. For example, the village map features mountainous terrain with tunnels, while the farmland map is characterized by diverse multicolored patterns. To overcome these limitations, this study utilizes Creo software [26] to model detailed scene elements, which are then integrated into the webots for constructing realistic maps.
Category Sensor Parameter Type Description Potential Tasks
Vision Camera Image Mat Images captured by the camera Default sensor
LiDAR LidarCloud vector2000 Point cloud from RPLIDAR A2 (m) Obstacle avoidance
Motion GPS Position vector3 Position (m) Visual navigation
Linear vector3 Linear velocity (m/s) Visual navigation
[…]
Gyroscope Angular vector3 Angular velocity (rad/s) Robot posture stabilization
IMU Angle vector3 Euler angles (rad) Robot posture stabilization
Orientation vector4 Quaternion representation Robot posture stabilization
Table 8: State parameters of DAT benchmark.
Parameter Type Description
cameraWidth double image width($px$)
cameraHeight double image height($px$)
cameraFov double camera field of view($rad$)
cameraF double estimated camera focal length($px$)
[…]
cameraMidGlobalPos vector3d coordinates of the camera center mapped to the ground in the world coordinate system
carMidGlobalPos vector3d coordinates of the vehicle center in the world coordinate system
cameraMidPos vector3d coordinates of the camera center in the world coordinate system
carDronePosOri vector4d 1D orientation + 3D position of vehicle in the drone coordinate system
crash double whether tracker collides with a building
carDir double car direction(0-stop,1-go straight,2-turn left,3-turn right)
carTypename string tracking target type
Table 9: Reward parameters of DAT benchmark. The homogeneous transformation matrices (HTM) $\texttt{T}_{\texttt{ct}}$ and $\texttt{T}_{\texttt{ct}}$ are 4$\times$4 square matrices. Therefore, their data type $\text{double}[16]$ corresponds to a double array of length 16.
Targets. All tracking target illustrations are presented in Fig. 6. Specifically, Fig. 6(a) presents automobile and motorbike tracking targets, including passenger vehicles (the first seven cars), buses, trucks, trailers, and motorcycles (such as scooters and motorbikes). These two categories of tracking targets leverage Simulation of Urban Mobility (SUMO) [30] for road behavior modeling and interaction management with other targets. In contrast, Fig. 6(b)-(d) display pedestrian, wheeled robot, and legged robot tracking targets, respectively. These three types of targets utilize SUMO paths for position initialization and rely on specific controllers for action and behavior management.
Sensors. In the real world VAT tasks, a single camera cannot ensure the agent’s stability and robustness. Thus, integration with other sensors is often required. The DAT benchmark provides common sensors that can obtain the drone’s state parameters relative to the world coordinate system. The drone’s position and velocity are determined using GPS, while its acceleration is measured by an accelerometer, providing essential self-referential data for visual navigation tasks. Angular velocity is recorded via a gyroscope, and Euler angles obtained from the IMU are converted into quaternions to facilitate state estimation and ensure orientation stability. Additionally, the RPLIDAR A2, provided by DAT, generates point cloud data, which supports tasks such as obstacle avoidance and navigation by delivering detailed environmental information. The specific sensors, their parameters, and potential tasks are summarized in Tab. 8.
Additional Parameters. The training process of VAT agents often requires additional parameters for effective reward design. To facilitate this, DAT benchmark provides 4 categories comprising a total of 13 parameters, supporting diverse reward design strategies, as detailed in Tab. 9.
First are the camera parameters, which mainly include image width cameraWidth, image height cameraHeight, field of view cameraFov, and focal length cameraF. Utilizing these, the camera plane can be projected onto the ground to aid in reward construction.
Next is the homogeneous transformation matrix (HTM). In the reward design, coordinate transformations are often required to express physical quantities within a unified coordinate system, enabling consistent calculations. For example, prior studies [18, 33, 17] transform the position, velocity, and acceleration of targets into the tracker’s coordinate system to construct rewards. To support such operations, DAT benchmark provides $\texttt{T}_{\texttt{ct}}$, the HTM mapping the drone camera coordinate system to the world coordinate system, and $\texttt{T}_{\texttt{tw}}$, the HTM mapping the tracking target’s coordinate system to the world coordinate system.
Additionally, for the state of the tracker itself, cameraMidPos represents the position of the drone camera’s optical center in the world coordinate system. The parameter crash indicates whether the drone collides with any buildings in the scene, which can be used in reward design for obstacle avoidance tasks.
Lastly, for ease of model training in simulations, reward design often depends on some privileged information, i.e., variables that are almost impossible to obtain in real-world settings. Thus, DAT benchmark also provides such adaptations. For example, carMidGlobalPos gives the target’s position in the world coordinate system, and carDronePosOri represents the target’s orientation and position relative to the drone coordinate system, frequently used in VAT reward design [18, 33, 17]. Furthermore, information on the target’s movement direction and type is provided, enabling training the drone’s predictive ability for target movement.
Figure 7: Schemetic diagram of reward acquisition under fixed-forward perspective. (a) Illustration of the mapping between the image plane $LU_{1}LD_{1}RD_{1}RU_{1}$ and the target plane $LU_{G1}LD_{G1}RD_{G1}RU_{G1}$, along with the tracker body-fixed frame $\{t\}$, whose $x$-axis is aligned with the forward flight direction of the drone. (b) Schematic of the reward acquisition plane. As this plane is rectangular, the design of the distance-based reward becomes effective.
Task Configuration. We encapsulate the scenes, tasks, and data augmentation into Python classes, and provide 3 different environment classes for different algorithm requirements. The base environment class directly interacts with webots and is designed to support asynchronous reinforcement learning algorithms, such as the asynchronous advantage actor-critic (A3C) algorithm [36]. The Gymnasium environment class wraps the base environment class into a Gymnasium [54] interface, enabling direct compatibility with popular reinforcement learning libraries, such as Stable-Baselines3 [43] and Tianshou [61] for efficient algorithm development and evaluation. The parallel environment class encapsulates the base environment class to enable parallel execution, providing direct support for synchronous algorithms, such as proximal policy optimization (PPO) [49] and soft actor-critic (SAC) [23]. Additionally, the scenario selection, tracker and target configuration, SUMO parameters, task additional parameters, and randomization methods can all be efficiently customized through a JSON configuration file.
Appendix C More Details of Proposed R-VAT
C.1 Theoretical Proof of Reward Design
Current reinforcement learning-based VAT methods assume a fixed forward-view perspective [73, 33, 17, 18] and design rewards based on physical distance. Distance-based reward design approaches fail to accurately reflect the performance of the agent in tracking tasks, especially in real-world applications that require a variable top-down perspective. In this section, we first demonstrate the effectiveness of distance-based reward design under the assumption of a fixed forward-view perspective. We then demonstrate that, under a top-down perspective, these approaches can result in scenarios where targets farther from the image center are assigned higher rewards than those closer to it.
The distance-based reward design approaches leverage the physical distance of the tracking target relative to the tracker to define the reward. While the exact reward formulations in existing methods differ, the underlying principle and the associated potential issues remain consistent. Therefore, we take the reward formulation from [18] as an example for the subsequent theoretical proof. The reward in [18] is expressed along the x, y, and z directions as follows:
$\small\begin{array}[]{rl}r_{x}(\cdot)\!&=\max{(0,1-\mid y_{x}(\cdot)-d_{r}\mid)},\\ r_{y}(\cdot)\!&=\max{(0,1-\mid\frac{2}{A_{FoV}}\arctan{(\frac{y_{y}(\cdot)}{y_{x}(\cdot)})}\mid)},\\ r_{z}(\cdot)\!&=\max{(0,1-\mid\frac{2}{A_{FoV}}\arctan{(\frac{y_{z}(\cdot)}{y_{x}(\cdot)})}\mid)},\end{array}$ (12)
where the reward is defined in the tracker body-fixed frame, as illustrated in Fig. 7(a).
Figure 8: Schematic diagram of pinhole imaging.
Effectiveness under a fixed forward-view perspective. As illustrated in Fig. 7(a), when the tracker adopts a fixed-forward perspective, the image plane is parallel to the plane containing the target. Based on the principle of pinhole imaging (see Fig. 8), the quadrilateral $LU_{1}LD_{1}RD_{1}RU_{1}$ on the image plane is geometrically similar to the quadrilateral $LU_{G1}LD_{G1}RD_{G1}RU_{G1}$ on the target plane, which forms a rectangle. Furthermore, the principle of pinhole imaging can also be utilized to establish the mapping range between the image plane and the target plane.
Figure 9: Schemetic diagram of reward acquisition under top-down perspective. (a) Illustration of the mapping between the image plane $LU_{2}LD_{2}RD_{2}RU_{2}$ and the target plane $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$. (b) Failure case of the distance-based reward design in the $x$ direction and the tracker body-fixed frame $\{t\}$, which is consistent with the definition in [18]. When the target is located in the trapezoid $abcd$, a target farther from the image center may receive a higher reward. (c) Failure case in the $y$ direction. When the target is located in triangular regions $a_{1}b_{1}c_{1}$ and $a_{2}b_{2}c_{2}$, points farther from the center point may receive a higher reward.
To simplify the analysis of the reward function shown in Eq. 12, we define a function $f(x)=\!1-\!\!\mid\!A\arctan x\!\mid$ where $A\!=\!\frac{2}{A_{FoV}}$ and $x\!=\!\frac{y_{y}(\cdot)}{y_{x}(\cdot)}$. The function is monotonically decreasing when $x\!\!>\!\!0$ and monotonically increasing when $x\!\!<\!\!0$. Specifically, given a fixed $y_{x}$, the relationship holds:
$\forall y_{y1}\!,y_{y2}\!\in\![-\frac{W}{2},\frac{W}{2}]\text{, if }\mid\!\!y_{y1}\!\!\mid>\mid\!\!y_{y2}\!\!\mid\!\text{, }f(\!\frac{y_{y1}}{y_{x}}\!)\!\!<\!\!f(\!\frac{y_{y2}}{y_{x}}\!),$ (13)
$\forall y_{z1}\!,y_{z2}\!\in\![-\frac{H}{2},\frac{H}{2}]\text{, if }\mid\!\!y_{z1}\!\!\mid>\mid\!\!y_{z2}\!\!\mid\!\text{, }f(\!\frac{y_{z1}}{y_{x}}\!)\!\!<\!\!f(\!\frac{y_{z2}}{y_{x}}\!).$
Consequently, across the entire image, the reward decreases as the distance of the target from the center of the image increases. This demonstrates that Eq. 12 provides an effective measure of drone tracking performance in both the $y$ and $z$ directions under the assumption of a fixed forward-view perspective (See Fig. 7(b)).
Figure 10: The zoomed-in illustration of the triangular region $I_{OG}RU_{G2}RD_{G2}$ shown in Fig. 9(c). The point $P_{T}$ represents the coordinate point of the tracker’s projection onto the ground.
Figure 11: Schematic diagram of the optical frustum in a camera. (a) The camera’s optical frustum $O_{c}\!-\!CRU_{G2}LU_{G2}D$ and its intersection with the ground, forming the quadrilateral $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$. This structure demonstrates the existence of the triangular region $a_{1}b_{1}c_{1}$ depicted in Fig. 9(c). (b) The zoomed-in illustration of the triangular regions $O_{c}EF$ and $O_{c}RU_{G2}LU_{G2}$, which are used in the theoretical proof presented in Eq. 17. (c) The zoomed-in illustration of the triangular regions $ORU_{G2}LU_{G2}$ and $NGH$, which are used in the theoretical proof presented in Eq. 18.
[…]
$\small S_{a^{\prime}bcd^{\prime}}\!=\!\{(x,y)\!\mid\!(x,y)\!\in\!\mathbb{R}^{2}\text{ and }(x,y)\text{ inside }a^{\prime}bcd^{\prime}\}.$ (14)
For any point $B\in S_{a^{\prime}bcd^{\prime}}$, it holds that $d(B)<d(A)$.
Moreover, according to Eq. 12, the reward $r_{x}(\cdot)$ in the $x$-direction is determined by $d_{x}(\cdot)=\mid y_{x}(\cdot)-d_{r}\mid$, which represents the distance in the x-direction of the target point from the image center. Based on this metric, the symmetric point $A^{\prime\prime}$ of point $A$ with respect to $I_{OG}$ and the iso-distance line segment $l_{ad}$ can be constructed. All points on $l_{ad}$ receive the same $x$-direction reward $r_{x}$ as point $A$.
Besides, the reward function $r_{x}$ in Eq. 12 decreases monotonically with increasing $d_{x}$. Thus, for any $d_{x}(B)>d_{x}(A)$, it follows that $r_{x}(B)<r_{x}(A)$. Accordingly, the set of points strictly inside trapezoid $abcd$ can be defined as:
$\small S_{abcd}\!=\!\{(x,y)\!\mid\!(x,y)\!\in\!\mathbb{R}^{2}\text{ and }(x,y)\text{ inside }abcd\}.$ (15)
For any $B\in S_{abcd}$, it follows that $r_{x}(B)<r_{x}(A)$, and $d(B)<d(A)$ as $B$ also lies within trapezoid $a^{\prime}bcd^{\prime}$. Therefore, it becomes evident that a point $A$, which is farther from the image center, may receive a higher reward than a point $B$, which is closer to the center. This highlights that the reward design in Eq. 12 fails to accurately represent tracking performance in the x-direction.
Failure Cases in the y-direction under Top-Down Perspective. In the $y$-direction, Fig. 9(c) illustrates another failure case of the reward, which occurs in the triangular regions $a_{1}b_{1}c_{1}$ and $a_{2}b_{2}c_{2}$. To ensure the validity of subsequent proofs, we first need to prove the existence of the triangular regions $a_{1}b_{1}c_{1}$ and $a_{2}b_{2}c_{2}$. This can be achieved by demonstrating that $\angle_{RU_{G2}OLU_{G2}}\!<\!FoV$ (See Sec. C.1 for theoretical proof), which ensures that the aforementioned triangular regions do not vanish due to parallel bounding edges. Subsequently, we analyze the failure case of the distance-based reward using the region $a_{1}b_{1}c_{1}$ as an example. A zoomed-in view is shown in Fig. 10.
As shown in Fig. 10, a target point $B$ is located within the region $a_{1}b_{1}c_{1}$, while another target point $A$ coincides with point $c_{1}$. According to the definition in Eq. 12, we can define a function $\theta(\cdot)=\arctan(\frac{y_{y}(\cdot)}{y_{x}(\cdot)})$, which represents the angle between the line segment $l_{P_{T}I_{OG}}$ and the line connecting the target coordinate point and the image center $I_{OG}$. For the target point $B$, $\theta(B)$ is defined as $\theta(B)=\angle_{BP_{T}I_{OG}}$, while for point $A$, $\theta(A)=\frac{FoV}{2}$. When the target point $B$ is within the region $a_{1}b_{1}c_{1}$, it holds that $\theta(B)<\theta(A)$. Furthermore, given that $\theta\!>\!0$, the function $r_{y}(\cdot)$ exhibits a monotonically decreasing behavior concerning $\theta$. As a result, it follows that $r_{y}(B)>r_{y}(A)$.
In addition, to determine the distance from point $B$ to the image center, a line segment $l_{BB^{\prime}}$ is constructed through point $B$ such that $l_{BB^{\prime}}\!\parallel\!l_{b_{1}A}$. Therefore, the inequality holds: $\frac{\mid l_{I_{OG}A}\mid}{\mid l_{I_{OG}RD_{G2}}\mid}\!<\!\frac{\mid l_{I_{OG}B^{\prime}}\mid}{\mid l_{I_{OG}RD_{G2}}\mid}$. Based on the definition of the distance from a point to the image center in Fig. 11, this implies that $d(B)>d(A)$. Consequently, the target point $B$, which is farther from the image center, receives a higher reward in the $y$-direction compared to the closer point $A$. This result demonstrates that the distance-based reward design fails to accurately reflect tracking performance in the $y$-direction.
Theoretical proof of angular relationships. To demonstrate that distance-based reward design is invalid in the top-down perspective (as discussed in Sec. C.1), it is first necessary to prove:
$\small\angle_{RU_{G2}OLU_{G2}}<FoV.$ (16)
The proof is based on the optical frustum shown in Fig. 11.
As shown in Fig. 11, the optical frustum $O_{c}\!-\!CRU_{G2}LU_{G2}D$ of the camera is depicted. The base quadrilateral $CRU_{G2}LU_{G2}D$ of the frustum is a rectangle. The line $l_{O_{c}O^{\prime}}$, which connects the apex $O_{c}$ to the ground center $O^{\prime}$, is perpendicular to the base plane. The intersection of the frustum with the ground plane produces the quadrilateral $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$. This quadrilateral represents the projection of the image plane onto the ground plane.
Thus, under the aforementioned notation, the proof of Eq. 16 can be divided into the following two steps. First, we demonstrate that when the point $RD_{G2}$ coincides with the apex $O_{c}$, the base angle of triangle $ORU_{G2}LU_{G2}$ is greater than that of triangle $O_{c}EF$, as follows:
$\small\angle_{ORU_{G2}LU_{G2}}>\angle_{O_{c}EF}.$ (17)
Next, we establish that as $RD_{G2}$ moves from point $C$ to point $O_{c}$ along the line segment $l_{CO_{c}}$, the base angle of triangle $ORU_{G2}LU_{G2}$ decreases monotonically. Formally, this can be expressed as:
$\small\forall RD_{G21},RD_{G22}\in l_{CO_{c}}\text{, if }\mid l_{CRD_{G21}}\!\mid>\mid l_{CRD_{G22}}\!\mid\text{,}\\ \text{ then }\angle_{ORU_{G21}LU_{G21}}\!<\!\angle_{ORU_{G22}LU_{G22}}$ (18)
As shown in Fig. 11(b), it is given that $\mid\!l_{EO^{{}^{\prime}}}\!\!\mid=\mid\!l_{RU_{G2}H}\!\!\mid$ and $\mid\!\!l_{O_{c}H}\!\!\mid>\mid\!\!l_{O_{c}O^{{}^{\prime}}}\!\!\!\mid$. Based on these conditions, the following inequality is satisfied: $\frac{\mid l_{O_{c}O^{{}^{\prime}}}\mid}{\mid l_{EO^{{}^{\prime}}}\mid}\!\!<\!\!\frac{\mid l_{O_{c}H}\mid}{\mid l_{RU_{G2}H}\mid}$. Moreover, the base angle of triangle $O_{c}RU_{G2}LU_{G2}$ and $O_{c}EF$ can be expressed as:
$\angle_{O_{c}EF}$ $=\arctan{(\frac{\mid l_{O_{c}O^{{}^{\prime}}}\mid}{\mid l_{EO^{{}^{\prime}}}\mid})},$ (19)
$\angle_{O_{c}RU_{G2}LU_{G2}}$ $=\arctan{(\frac{\mid l_{O_{c}H}\mid}{\mid l_{RU_{G2}H}\mid})}.$
Given that the function $\arctan(\cdot)$ is monotonically increasing, it follows that: $\angle_{O_{c}RU_{G2}LU_{G2}}\!>\!\angle_{O_{c}EF}$.
As shown in Fig. 11(c), a line $l_{RD_{G2}R^{\prime}}$ is constructed through point $RD_{G2}$, perpendicular to $l_{RU_{G2}LU_{G2}}$. The length of this perpendicular segment is equal to the height of the trapezoid, i.e., $\mid l_{NH}\mid=\mid l_{RD_{G2}R^{{}^{\prime}}}\mid$. Consequently, the base angle of $ORU_{G2}LU_{G2}$ can be expressed as:
$\small\angle_{ORU_{G2}LU_{G2}}=\arctan{(\frac{\mid l_{NH}\mid}{\mid l_{RU_{G2}R^{{}^{\prime}}}\mid})}.$ (20)
In the trapezoid $LU_{G2}LD_{G2}RD_{G2}RU_{G2}$, the length of line segment $l_{RU_{G2}R^{{}^{\prime}}}$ is given by the following equation:
$\small\mid\!l_{RU_{G2}R^{{}^{\prime}}}\!\mid=\frac{1}{2}(\mid\!l_{RU_{G2}LU_{G2}}\!\mid\!-\!\mid\!l_{RD_{G2}LD_{G2}}\!\mid).$ (21)
Additionally, in triangle $O_{c}CD$, the relationship exists:
$\small\frac{\mid l_{RD_{G2}LD_{G2}}\mid}{\mid l_{CD}\mid}=1-\frac{\mid l_{NG}\mid}{\mid l_{O_{c}G}\mid}.$ (22)
By combining Eq. 21 and Eq. 22, and given that $\mid\!l_{RU_{G2}LU_{G2}}\!\mid=\mid\!l_{CD}\!\mid$ holds in rectangle $CRU_{G2}LU_{G2}D$, the length of line segment $l_{RU_{G2}R^{{}^{\prime}}}$ can be determined as:
$\small\mid l_{RU_{G2}R^{{}^{\prime}}}\mid=\frac{1}{2}\frac{\mid l_{NG}\mid\cdot\mid l_{CD}\mid}{\mid l_{O_{c}G}\mid}.$ (23)
In triangle $NGH$, by applying the law of cosines, the length of line segment $l_{NH}$ can be calculated as:
$\small\mid\!l_{NH}\!\mid=\sqrt{\mid\!l_{NG}\!\mid^{2}\!+\!\mid\!l_{GH}\!\mid^{2}\!-2\mid\!l_{NG}\!\mid\mid\!l_{GH}\!\mid\!\cos\theta}.$ (24)
By combining Eq. 23 and Eq. 24, the following relationship can be derived:
$\small\frac{\mid l_{NH}\mid}{\mid l_{RU_{G2}R^{{}^{\prime}}}\mid}=c\sqrt{1+x^{2}-2x\cos\theta},$ (25)
where $c=\frac{2\mid l_{O_{c}G}\mid}{\mid l_{CD}\mid}$ and $x=\frac{\mid l_{GH}\mid}{\mid l_{NG}\mid}$. Furthermore, in the right triangle $O_{c}O^{{}^{\prime}}G$, the value of $\cos\theta$ is given by: $\cos\theta=\frac{\mid l_{GH}\mid}{2\mid l_{O_{c}G}\mid}$. Since point $RD_{G2}$ moves along segment $l_{CO_{c}}$, it follows that $\mid l_{NG}\mid\in[0,\mid l_{O_{c}G}\mid]$, which implies:
$\small x>\cos\theta.$ (26)
We can further define function $g(x)=x^{2}\!-\!2x\cos\theta\!+\!1$ and its derivative $g^{{}^{\prime}}(x)=2(x-\cos\theta)$. According to Eq. 26, $g^{{}^{\prime}}(x)>0$, which indicates that $g(x)$ is monotonically increasing with respect to $x$.
Additionally, since $h(\cdot)=c\sqrt{\cdot}$ and $\arctan(\cdot)$ are monotonically increasing, and $x=\frac{\mid l_{GH}\mid}{\mid l_{NG}\mid}$ is monotonically decreasing with respect to $\mid l_{NG}\mid$, the monotonicity rule for composite functions ensures that $\angle_{ORU_{G2}LU_{G2}}$ in Eq. 20 is monotonically decreasing regarding $\mid l_{NG}\mid$. Therefore, the minimum value of $\angle_{ORU_{G2}LU_{G2}}$ is achieved when $\mid\!l_{NG}\!\mid$ reaches its maximum value, which equals $\mid\!l_{O_{c}G}\!\mid$, i.e., when $RD_{G2}$ coincides with $O_{c}$. The minimum value is given by:
$\small\min\{\angle_{ORU_{G2}LU_{G2}}\}=\angle_{O_{c}RU_{G2}LU_{G2}},$ (27)
which implies that: $\angle_{RU_{G2}OLU_{G2}}<FoV$.
Figure 12: Network structure of Drone Agent.
C.2 More Details
[…]
Data Augmentation. In our two-stage curriculum learning process, we employ identical data augmentation. The flight altitude is selected from the interval $[13,22]\text{m}$, and the camera pitch angle is chosen from $[0.6,1.38]\text{rad}$. These parameters are consistent throughout each episode. Meanwhile, the drone’s initial orientation relative to the target fluctuates within the range $[-\pi,\pi]\text{rad}$, and the target’s initial position is set between $[-4.5,-2.5]\cup[2.5,4.5]\text{m}$.
Figure 13: Schematic diagram of the training environments for the two-stage of Curriculum Learning
Details on coordinate transformations. Given two planes $P_{0}:\hat{n_{0}}\textbf{x}^{T}+D_{0}=0$ and $P_{1}:\hat{n_{1}}\textbf{x}^{T}+D_{1}=0$, along with the HTM $T_{01}$ from $P_{0}$ to $P_{1}$. The transformation $T_{01}$ is defined as
$\small T_{01}=\begin{bmatrix}R_{01}&t_{01}\\ 0&1\\ \end{bmatrix}.$ (28)
Hence, the expression of plane $P_{1}$ can be obtained using the analytical expression of plane $P_{0}$ and $T_{01}$ as follows:
$\hat{n_{1}}^{T}$ $=R_{01}\hat{n_{0}}^{T},$ (29)
$D_{1}$ $=D_{0}-\hat{n_{1}}t_{01}.$
[…]
$\small T_{ct}=\begin{bmatrix}R_{ct}&t_{ct}\\ 0&1\\ \end{bmatrix},$ (30)
[…]
$\small T_{tc}=\begin{bmatrix}R^{T}_{ct}&-R^{T}_{ct}t_{ct}\\ 0&1\\ \end{bmatrix}.$ (31)
Using Eq. 29 and the transformation matrix $T_{tc}$, the plane $G_{c}$ can be formulated as $P_{G_{c}}=(r_{3}^{T},-h+r_{3}^{T}R^{T}_{ct}t_{ct})$.
Sparse Reward. In addition to the dense reward function described in the main text, we also provide a sparse reward function design. The sparse reward only provides a fixed reward when the target is within the image and no reward when it is outside. The definition of $r_{d}$ is as follows.
$\small r_{d}=\left\{\begin{array}[]{lr}1,&t\in I\\ 0,&\text{otherwise},\end{array}\right.$ (32)
where $I$ represents the image range of the drone. This reward can be used to construct the experimental metric, Tracking Success Rate (TSR).
Training algorithm. For the training method of R-VAT, we choose to use PPO algorithm. PPO algorithm regulates the speed of gradient updates by constraining the magnitude of policy changes $r_{t}$, expressed as follows:
$\small r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{old}}(a_{t}|s_{t})},$ (33)
where $\pi_{\theta}$ and $\pi_{\theta_{old}}$ are the new and old policies. Additionally, to enhance the agent’s exploration, we introduce an entropy loss term $\mathcal{H}$, formulated as:
$\small\mathcal{H}(\pi_{\theta}(s))=-\sum_{a}\pi_{\theta}(a|s)\log{\pi_{\theta}(a|s)}.$ (34)
The optimization objective for the actor is as follows:
$\small\mathcal{L}_{A}=\hat{\mathbb{E}}[\min{(r_{t}\hat{A_{t}},\text{clip}(r_{t},1\!-\!\epsilon,1\!+\!\epsilon)\hat{A_{t}})}\!+\!\beta\mathcal{H}],$ (35)
where $\hat{A_{t}}$ is the advantage function, $\epsilon$ is the clip parameter, and $\beta$ is the entropy coefficient. The expression of $\hat{A_{t}}$ is computed as:
$\small\hat{A_{t}}=\sum_{l=0}^{E_{l}-t}{(\gamma\lambda)^{l}\delta_{t+l}},$ (36)
where $T,\lambda,\delta_{t+l}$ are the data collection step, generalized advantage estimator (GAE) [50] discount factor and temporal difference error respectively. The optimization objective expression of the critic network $V$ is defined as:
$\small\mathcal{L}_{C}=\hat{\mathbb{E}}_{t}[(r_{t}+\gamma V(s_{t+1})-V(s_{t}))^{2}].$ (37)
The hyperparameters of the PPO algorithm used in this article are set as follows: discount factor $\gamma=0.9$, GAE discount factor $\lambda=0.95$, entropy coefficient $\beta=0.01$, PPO clipping parameter $\epsilon=0.2$.
Curriculum Learning for Agent Training. We introduce a Curriculum-Based Training (CBT) strategy designed to progressively enhance the performance of the tracker. In the first-stage curriculum, the agent is trained to track vehicles moving along straight trajectories without occlusions or extra interference. In the second-stage curriculum, the agent is exposed to visually complex environments and tasked with tracking targets exhibiting diverse and dynamic behaviors. The scenario of each stage is shown in Fig. 13, where the upper row is the first-stage environment, and the lower row corresponds to the second-stage environment.
Appendix D Baselines
Active Object Tracking (AOT) [33]. In this paper, the agent learns to follow a fixed target-tracking trajectory using the A3C algorithm. In addition, the Agent uses the following reward function:
$\small r=A-(\frac{\sqrt{x^{2}+(y-d)^{2}}}{c}+\lambda\mid\omega\mid),$ (38)
where $d$ represents the optimal distance between the tracker and the target, $c$ is the maximum allowable distance between the tracker and the target, and $A$ denotes the maximum reward. In the original paper, $c=200$ and $A=1.0$. During our replication, we set $A=1.0$, but due to the drone’s camera being tilted downward, a value of $c=200$ would far exceed the camera’s field of view, which is unrealistic. Therefore, we modify the parameter $c$ to be the maximum offset distance that keeps the target within the image frame, i.e., $c=9$.
Scene citystreet desert village downtown lake farmland
Steps (M) 19.2 13.4 21.3 19.8 9.9 9.2
Table 10: Total training steps on different scenes. During the training process, we employ a parallel training approach involving 35 agents. Consequently, the reported total training steps represent the cumulative steps taken by all agents combined.
Scene citystreet desert village downtown lake farmland
T (M) 10.0 6.2 8.0 10.3 5.6 4.1
Table 11: Transition steps across different scenes.
D-VAT[18]. In this approach, the agent uses an asymmetric Actor-Critic network structure and the soft actor-critic learning method [23] to accomplish the task of drone tracking another drone. In the actual comparative experiments, we convert it from a continuous action space to a discrete action space, referring to [10]. Additionally, the method uses the following reward function.
$\small r(k)=\left\{\!\!\begin{array}[]{l c}\!\!r_{e}(k)\!-\!k_{v}r_{v}(k)\!-\!k_{u}r_{u}(k)&\!\|y(k)\|\!>\!d_{m}\\[2.84526pt] \!\!-k_{c}&\text{otherwise},\end{array}\right.$ (39)
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
AOT $49_{\pm 3}$ $49_{\pm 9}$ $45_{\pm 5}$ $49_{\pm 3}$ $48_{\pm 3}$ $48_{\pm 3}$ $49_{\pm 4}$ $49_{\pm 3}$ $49_{\pm 3}$
D-VAT $48_{\pm 8}$ $46_{\pm 12}$ $46_{\pm 10}$ $57_{\pm 11}$ $50_{\pm 8}$ $46_{\pm 3}$ $48_{\pm 9}$ $54_{\pm 10}$ $53_{\pm 10}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
AOT $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 0}$ $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$
D-VAT $51_{\pm 10}$ $47_{\pm 13}$ $46_{\pm 10}$ $56_{\pm 11}$ $39_{\pm 8}$ $47_{\pm 3}$ $48_{\pm 13}$ $48_{\pm 13}$ $39_{\pm 10}$
R-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $\textbf{119}_{\pm 110}$ $\textbf{170}_{\pm 139}$ $\textbf{275}_{\pm 121}$ $\textbf{182}_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
AOT $51_{\pm 7}$ $51_{\pm 11}$ $46_{\pm 5}$ $49_{\pm 4}$ $52_{\pm 11}$ $57_{\pm 24}$ $47_{\pm 5}$ $47_{\pm 5}$ $47_{\pm 5}$
D-VAT $46_{\pm 8}$ $45_{\pm 9}$ $44_{\pm 8}$ $69_{\pm 42}$ $45_{\pm 8}$ $45_{\pm 3}$ $44_{\pm 8}$ $44_{\pm 8}$ $43_{\pm 8}$
R-VAT $\textbf{234}_{\pm 122}$ $\textbf{160}_{\pm 139}$ $\textbf{239}_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $\textbf{140}_{\pm 118}$ $\textbf{257}_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
AOT $52_{\pm 3}$ $52_{\pm 9}$ $48_{\pm 7}$ $54_{\pm 5}$ $53_{\pm 5}$ $54_{\pm 8}$ $54_{\pm 5}$ $54_{\pm 5}$ $54_{\pm 5}$
D-VAT $8_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 2}$
R-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
AOT $49_{\pm 3}$ $49_{\pm 10}$ $46_{\pm 5}$ $49_{\pm 3}$ $47_{\pm 3}$ $49_{\pm 3}$ $48_{\pm 3}$ $48_{\pm 4}$ $48_{\pm 3}$
D-VAT $50_{\pm 8}$ $45_{\pm 9}$ $45_{\pm 10}$ $70_{\pm 42}$ $46_{\pm 8}$ $43_{\pm 2}$ $46_{\pm 8}$ $51_{\pm 8}$ $49_{\pm 9}$
R-VAT $\textbf{112}_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $\textbf{181}_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $\textbf{168}_{\pm 110}$ $\textbf{99}_{\pm 67}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
AOT $51_{\pm 7}$ $50_{\pm 9}$ $46_{\pm 5}$ $49_{\pm 3}$ $51_{\pm 9}$ $60_{\pm 25}$ $48_{\pm 4}$ $56_{\pm 24}$ $56_{\pm 24}$
D-VAT $13_{\pm 2}$ $13_{\pm 1}$ $13_{\pm 1}$ $15_{\pm 1}$ $14_{\pm 1}$ $13_{\pm 1}$ $14_{\pm 1}$ $13_{\pm 1}$ $14_{\pm 1}$
R-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 12: The detailed results of comparison experiments on $CR$ metric. Each row in the table represents the training results on a specific map, e.g., the first row corresponds to the results trained on the map citystreet-day. Each column represents the testing results on the corresponding map, e.g., the second column corresponds to the testing results on the map desert-day.
Figure 14: Schematic diagram of reward curves on DAT scenes. The red curve represents the reward of the proposed R-VAT, while the yellow curve corresponds to the R-VAT (w/o CBT).
In the above equation Eq. 39, $r_{v}(k)$ and $r_{u}(k)$ are regularization terms for the drone’s speed and output control, as shown in Eq. 40. For the discrete action space, the regularization term has a fixed value for a given action. This term only regularizes the linear velocity of the drone, which causes the drone to tend to perform rotational movements. Therefore, in the reproduction process, we set $k_{v}=0$ and $k_{u}=0$. Additionally, due to the unexpectedly large acceleration values obtained for the target relative to the tracker under the discrete action setting, we set the input acceleration of the critic network to $a(k)=0$.
$\small r_{v}(k)=\dfrac{\|v(k)\|}{1+\|v(k)\|},\;\;r_{u}(k)=\dfrac{\|u(k)\|}{1+\|u(k)\|}.$ (40)
It is important to note that in the AOT and D-VAT experiments, the target is initially positioned at the center of the tracker’s image, and the initial forward directions of both the tracker and the target are aligned. Additionally, since the success criterion of DAT requires the agent to keep the target at the center of its view, the optimal distance between the tracker and the target is defined as the distance in the forward direction when the target is at the center of the camera’s field of view. The tracker’s flight altitude is set to 22 meters, and the gimbal pitch angle is 1.37 radians, which remains consistent with the parameters used during testing.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.25_{\pm 0.02}$ $0.24_{\pm 0.03}$ $0.22_{\pm 0.03}$ $0.25_{\pm 0.02}$ $0.23_{\pm 0.03}$ $0.24_{\pm 0.01}$ $0.25_{\pm 0.02}$ $0.25_{\pm 0.02}$ $0.24_{\pm 0.02}$
D-VAT $0.26_{\pm 0.02}$ $0.25_{\pm 0.04}$ $0.25_{\pm 0.02}$ $0.32_{\pm 0.08}$ $0.27_{\pm 0.04}$ $0.19_{\pm 0.01}$ $0.26_{\pm 0.02}$ $0.28_{\pm 0.02}$ $0.29_{\pm 0.02}$
R-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.01}$
D-VAT $0.27_{\pm 0.02}$ $0.26_{\pm 0.04}$ $0.25_{\pm 0.02}$ $0.32_{\pm 0.07}$ $0.23_{\pm 0.03}$ $0.26_{\pm 0.01}$ $0.26_{\pm 0.04}$ $0.26_{\pm 0.04}$ $0.26_{\pm 0.04}$
R-VAT $\textbf{0.73}_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $\textbf{0.38}_{\pm 0.32}$ $\textbf{0.56}_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $\textbf{0.57}_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.25_{\pm 0.03}$ $0.25_{\pm 0.04}$ $0.23_{\pm 0.03}$ $0.24_{\pm 0.02}$ $0.25_{\pm 0.02}$ $0.26_{\pm 0.06}$ $0.23_{\pm 0.03}$ $0.23_{\pm 0.03}$ $0.23_{\pm 0.03}$
D-VAT $0.23_{\pm 0.04}$ $0.23_{\pm 0.04}$ $0.22_{\pm 0.05}$ $0.31_{\pm 0.14}$ $0.24_{\pm 0.06}$ $0.22_{\pm 0.01}$ $0.22_{\pm 0.04}$ $0.22_{\pm 0.05}$ $0.23_{\pm 0.05}$
R-VAT $\textbf{0.72}_{\pm 0.28}$ $\textbf{0.51}_{\pm 0.34}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $\textbf{0.48}_{\pm 0.31}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.40}_{\pm 0.29}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.30_{\pm 0.04}$ $0.26_{\pm 0.05}$ $0.27_{\pm 0.02}$ $0.29_{\pm 0.01}$ $0.29_{\pm 0.03}$ $0.29_{\pm 0.02}$ $0.29_{\pm 0.01}$ $0.29_{\pm 0.01}$ $0.29_{\pm 0.01}$
D-VAT $0.05_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.25_{\pm 0.02}$ $0.25_{\pm 0.03}$ $0.23_{\pm 0.03}$ $0.24_{\pm 0.02}$ $0.24_{\pm 0.02}$ $0.24_{\pm 0.01}$ $0.24_{\pm 0.01}$ $0.24_{\pm 0.02}$ $0.24_{\pm 0.01}$
D-VAT $0.25_{\pm 0.04}$ $0.23_{\pm 0.04}$ $0.23_{\pm 0.05}$ $0.30_{\pm 0.15}$ $0.26_{\pm 0.06}$ $0.22_{\pm 0.01}$ $0.26_{\pm 0.06}$ $0.26_{\pm 0.06}$ $0.25_{\pm 0.06}$
R-VAT $\textbf{0.43}_{\pm 0.25}$ $\textbf{0.47}_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $\textbf{0.43}_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $\textbf{0.59}_{\pm 0.39}$ $\textbf{0.62}_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.24_{\pm 0.02}$ $0.24_{\pm 0.04}$ $0.22_{\pm 0.03}$ $0.25_{\pm 0.02}$ $0.24_{\pm 0.02}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$
D-VAT $0.07_{\pm 0.01}$ $0.07_{\pm 0.01}$ $0.07_{\pm 0.00}$ $0.08_{\pm 0.01}$ $0.07_{\pm 0.00}$ $0.07_{\pm 0.00}$ $0.08_{\pm 0.00}$ $0.07_{\pm 0.00}$ $0.08_{\pm 0.00}$
R-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Table 13: The detailed results of comparison experiments on $TSR$ metric. Each row in the table represents the training results on a specific map, e.g., the second row corresponds to the results trained on the map desert-day. Each column represents the testing results on the corresponding map, e.g., the second column corresponds to the testing results on the map desert-day.
Appendix E More Experiments
E.1 Settings
Training Steps. Due to the varying challenges posed by different scene maps, the convergence speed of the agent differs across experiments, resulting in a unique total training step count for each map. The number of training steps for the agent across the six scenarios is shown in Tab. 10.
Ablation Experiment Settings. In this section, we introduce the training conditions of the single-stage reinforcement learning method and R-VAT, as well as the criteria for stage transitions. In single-stage reinforcement learning, the agent is placed in one of six scenarios (citystreet, desert, village, downtown, lake, and farmland) for training. For R-VAT, the agent is first trained in an environment where a randomly colored target moves straight along a line without obstacles. After convergence, the model is then trained in the corresponding complex scenarios. Specifically, the transition steps T for R-VAT are shown in Tab. 11.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $54_{\pm 7}$ $37_{\pm 21}$ $30_{\pm 6}$ $30_{\pm 14}$ $48_{\pm 13}$ $48_{\pm 4}$ $54_{\pm 9}$ $54_{\pm 9}$ $54_{\pm 9}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $253_{\pm 132}$ $302_{\pm 99}$ $284_{\pm 92}$ $\textbf{175}_{\pm 102}$ $\textbf{236}_{\pm 123}$ $266_{\pm 110}$ $\textbf{241}_{\pm 127}$ $279_{\pm 120}$ $306_{\pm 95}$
R-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $119_{\pm 110}$ $170_{\pm 139}$ $\textbf{275}_{\pm 121}$ $182_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $230_{\pm 120}$ $\textbf{197}_{\pm 124}$ $\textbf{255}_{\pm 118}$ $59_{\pm 69}$ $126_{\pm 105}$ $\textbf{182}_{\pm 120}$ $\textbf{267}_{\pm 93}$ $208_{\pm 141}$ $73_{\pm 68}$
R-VAT $\textbf{234}_{\pm 122}$ $160_{\pm 139}$ $239_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $140_{\pm 118}$ $257_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $54_{\pm 9}$ $49_{\pm 13}$ $47_{\pm 8}$ $57_{\pm 15}$ $51_{\pm 9}$ $48_{\pm 4}$ $29_{\pm 3}$ $57_{\pm 15}$ $58_{\pm 15}$
R-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $\textbf{124}_{\pm 90}$ $88_{\pm 52}$ $191_{\pm 108}$ $93_{\pm 75}$ $\textbf{187}_{\pm 123}$ $198_{\pm 117}$ $183_{\pm 110}$ $\textbf{185}_{\pm 102}$ $\textbf{102}_{\pm 57}$
R-VAT $112_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $181_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $168_{\pm 110}$ $99_{\pm 67}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $52_{\pm 9}$ $47_{\pm 9}$ $45_{\pm 9}$ $69_{\pm 42}$ $50_{\pm 9}$ $46_{\pm 2}$ $46_{\pm 2}$ $46_{\pm 3}$ $46_{\pm 2}$
R-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 14: Effectiveness of Curriculum-Based Training strategy on the DAT benchmark, results from $CR$ metric.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $0.30_{\pm 0.05}$ $0.14_{\pm 0.10}$ $0.20_{\pm 0.10}$ $0.31_{\pm 0.15}$ $0.28_{\pm 0.06}$ $0.21_{\pm 0.01}$ $0.30_{\pm 0.05}$ $0.30_{\pm 0.05}$ $0.30_{\pm 0.05}$
R-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $\textbf{0.83}_{\pm 0.28}$ $0.75_{\pm 0.32}$ $0.66_{\pm 0.34}$ $\textbf{0.52}_{\pm 0.28}$ $\textbf{0.69}_{\pm 0.24}$ $0.74_{\pm 0.26}$ $\textbf{0.59}_{\pm 0.36}$ $0.74_{\pm 0.34}$ $0.75_{\pm 0.34}$
R-VAT $0.73_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $0.38_{\pm 0.32}$ $0.56_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $0.57_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $\textbf{0.73}_{\pm 0.28}$ $\textbf{0.62}_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.16}$ $0.23_{\pm 0.17}$ $0.46_{\pm 0.25}$ $\textbf{0.58}_{\pm 0.33}$ $\textbf{0.71}_{\pm 0.28}$ $0.69_{\pm 0.33}$ $\textbf{0.40}_{\pm 0.24}$
R-VAT $0.72_{\pm 0.28}$ $0.51_{\pm 0.34}$ $0.73_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $0.48_{\pm 0.31}$ $0.71_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $0.40_{\pm 0.29}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $0.29_{\pm 0.04}$ $0.27_{\pm 0.03}$ $0.27_{\pm 0.03}$ $0.33_{\pm 0.06}$ $0.28_{\pm 0.03}$ $0.27_{\pm 0.01}$ $0.33_{\pm 0.06}$ $0.33_{\pm 0.06}$ $0.33_{\pm 0.06}$
R-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $\textbf{0.51}_{\pm 0.30}$ $\textbf{0.47}_{\pm 0.29}$ $0.45_{\pm 0.22}$ $\textbf{0.44}_{\pm 0.23}$ $0.57_{\pm 0.28}$ $0.59_{\pm 0.26}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.62}_{\pm 0.24}$ $0.33_{\pm 0.15}$
R-VAT $0.43_{\pm 0.25}$ $0.47_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $0.43_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $0.59_{\pm 0.39}$ $0.62_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $0.26_{\pm 0.04}$ $0.24_{\pm 0.04}$ $0.23_{\pm 0.05}$ $0.31_{\pm 0.14}$ $0.26_{\pm 0.06}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$
R-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Table 15: Effectiveness of Curriculum-Based Training strategy on the DAT benchmark, results from $TSR$ metric.
E.2 Comparison Experiments
We provide a comprehensive analysis of the comparative experimental results. Specifically, we provide detailed evaluations for within-scene testing (same scenes, same weather), cross-scene testing (same weather, different scenes) and cross-domain testing (same scene, different weather). Tab. 12 reports the $CR$ metric of three models under cross-scene and cross-domain conditions, while Tab. 13 presents the $TSR$ metric.
As shown in Tab. 12 and Tab. 13, the proposed R-VAT significantly outperforms existing state-of-the-art methods. Due to the reward design based on physical distance, both the AOT [33] and D-VAT [18] fail to accurately reflect the agent’s tracking performance from a top-down perspective (see Sec. C.1 for theoretical proof), leading to misleading training signals for the tracker. Consequently, neither AOT nor D-VAT can effectively learn meaningful features, resulting in irregular performance distributions. In contrast, the proposed R-VAT achieves superior convergence across all scenes. Specifically, in cross-scene experiments, the testing performance of the agent on the downtown map is relatively low, indicating that dense buildings and complex road elements pose significant challenges to the agent. Conversely, the testing performance on the village map is comparatively high, suggesting that the uniform color and simpler road conditions in the village map present fewer challenges.
For cross-domain testing experiments, the agent performs well under night and foggy conditions but struggles under snow conditions. This indicates that the proposed R-VAT exhibits strong robustness to changes in lighting and visibility but is less adaptive to variations in scene tone.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $9_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 0}$ $8_{\pm 1}$ $9_{\pm 0}$ $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $9_{\pm 1}$ $9_{\pm 0}$ $8_{\pm 1}$ $9_{\pm 0}$ $8_{\pm 0}$ $10_{\pm 0}$ $8_{\pm 1}$ $10_{\pm 1}$ $8_{\pm 0}$
R-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $\textbf{119}_{\pm 110}$ $\textbf{170}_{\pm 139}$ $\textbf{275}_{\pm 121}$ $\textbf{182}_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $9_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 0}$ $8_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$
R-VAT $\textbf{234}_{\pm 122}$ $\textbf{160}_{\pm 139}$ $\textbf{239}_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $\textbf{140}_{\pm 118}$ $\textbf{257}_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $8_{\pm 1}$ $8_{\pm 0}$ $8_{\pm 1}$ $9_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 0}$
R-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $11_{\pm 3}$ $11_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 2}$ $9_{\pm 0}$ $8_{\pm 0}$ $9_{\pm 0}$ $10_{\pm 1}$ $8_{\pm 1}$
R-VAT $\textbf{112}_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $\textbf{181}_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $\textbf{168}_{\pm 110}$ $\textbf{99}_{\pm 67}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $9_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 0}$ $9_{\pm 0}$ $9_{\pm 0}$
R-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 16: Effectiveness of reward design on the DAT benchmark, results from $CR$ metric.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.10_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.09_{\pm 0.01}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.73}_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $\textbf{0.38}_{\pm 0.32}$ $0.56_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $\textbf{0.57}_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.72}_{\pm 0.28}$ $\textbf{0.51}_{\pm 0.34}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $\textbf{0.48}_{\pm 0.31}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.40}_{\pm 0.29}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.10_{\pm 0.01}$ $0.09_{\pm 0.01}$ $0.07_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.08_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.43}_{\pm 0.25}$ $\textbf{0.47}_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $\textbf{0.43}_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $\textbf{0.59}_{\pm 0.39}$ $\textbf{0.62}_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Table 17: Effectiveness of reward design on the DAT benchmark, results from $CR$ metric. Each row in the table represents the training results on a specific map, e.g., the first row corresponds to the results trained on the map citystreet-day. Each column represents the testing results on the corresponding map, e.g., the second column corresponds to the testing results on the map desert-day.
E.3 Ablation Experiments
We present a comprehensive analysis of the ablation studies. First, we provide the reward curves for the citystreet, desert, village, downtown, lake, and farmland maps (see Fig. 14). Next, we provide detailed experimental results on the effectiveness of the Curriculum-Based Training strategy, as shown in Tab. 14 and Tab. 15.
Finally, the effectiveness of the reward in the R-VAT can be found in Tab. 16 and Tab. 17.
Effectiveness of Curriculum-Based Training strategy. To validate the effectiveness of the proposed Curriculum-Based Training (CBT) strategy, we conduct ablation experiments by removing the CBT module. The results for the $CR$ and $TSR$ metrics are presented in Tab. 14 and Tab. 15, respectively. The experimental results demonstrate that single-stage reinforcement learning methods without the CBT strategy successfully learn task objectives and achieve convergence on the desert, village, and lake maps. These three maps exhibit similar environmental characteristics: the desert and village maps feature uniform background colors and relatively simple road elements. Although the desert map has road segments partially covered by sand, these challenges are easy for the agent to overcome. Similarly, while the village map includes tunnels that may block vision, the proportion of tunnels is low. Additionally, although the lake map exhibits diverse background colors, the diversity primarily arises from vegetation-covered areas, which occupy a small proportion of the map, resulting in low challenges for the agent. In contrast, single-stage reinforcement learning methods without the CBT strategy fail to converge on the citystreet, downtown, and farmland maps. This suggests that as the visual complexity of scenes and the density of elements increase, directly applying single-stage reinforcement learning is highly challenging and unlikely to converge. These results demonstrate the effectiveness of the CBT strategy.
Effectiveness of reward design. To experimentally validate the effectiveness of the reward design proposed in this paper and to corroborate the theoretical proof in Sec. C.1, we conduct ablation experiments on the reward function. The comparative method utilizes the reward function from [18]. The detailed experimental results for the $CR$ and $TSR$ metrics are provided in Tab. 16 and Tab. 17. For within-scene testing, the R-VAT achieves an average improvement of $1100\%(0.06\rightarrow 0.72)$ in the $TSR$ metric compared to the reward design in [18]. In cross-scene and cross-domain testing, the R-VAT achieves average enhancements of $850\%(0.06\rightarrow 0.57)$ and $1017\%(0.06\rightarrow 0.67)$ in the $TSR$ metric, respectively. These results strongly demonstrate the high effectiveness of the proposed reward.